\begin{thebibliography}{99}
	
	
	\vspace*{1mm}
	\footnotesize
\bibitem{Sacks_vocal_tract}T. Aktosun,  A. Machuca and P. Sacks;	Determining the shape of a human vocal tract from pressure measurements at the lips, Inverse Problems, vol. 33, (2017), 115002, 33 pages.
		\bibitem{Bube and Burridge 1-D}K.P.  Bube and R. Burridge;
	The one-dimensional inverse problem of reflection seismology.
	SIAM Rev. 25 (1983), no. 4, 497--559.
		\bibitem{Burridge integral equation}R. Burridge; The Gel'fand-Levitan, the Marchenko, and the Gopinath-Sondhi integral equations of inverse scattering theory, regarded in the context of inverse impulse-response problems. Wave Motion 2 (1980), no. 4, 305--323.
	\bibitem{Eemeli}E. Bl\r{a}sten;	Well-posedness of the Goursat problem and stability for point source inverse backscattering, 2017 Inverse Problems 33 125003.
		\bibitem{Friedlander book}F.G.  Friedlander;  The wave equation on a curved space-time. Cambridge University Press, Cambridge-New York-Melbourne, 1975. Cambridge Monographs on Mathematical Physics, No. 2.
	\bibitem{Friedlander-Distributions-book} F.G.  Friedlander and M. Joshi; The Theory of Distributions, 2nd Edition, Cambridge University Press, 1998.

	\bibitem{Gelfand} I.M.  Gel'fand and G. E. Shilov;  Generalized functions. Vol. 1. Properties and operations. Translated from the 1958 Russian original by Eugene Saletan. 
	
 	\bibitem{Levrentev Romanov and Shishatski book ill-posed}M. M. Lavrent'ev,  V. G. Romanov and  S. P. Shishat$\cdot$ski\~{i}; Ill-posed problems of mathematical physics and analysis. Translated from the Russian by J. R. Schulenberger. 
 \bibitem{Ning_Yamamoto}W. Ning and M. Yamamoto;	The Gel’fand–Levitan theory for one-dimensional
 	hyperbolic systems with impulsive inputs, Inverse Problems 24 (2008) 025004 (19pp).
 	\bibitem{Rakesh impedance inversion} Rakesh; An inverse impedance transmission problem for the wave equation. Comm. Partial Differential Equations 18 (1993), no. 3-4, 583--600. 
 	\bibitem{Rakesh and Sacks impedance inversion} Rakesh and P. Sacks; Impedance inversion from transmission data for the wave equation. Wave Motion 24 (1996), no. 3, 263--274. 
 	\bibitem{Rakesh inversion of spherically symmetric potential} Rakesh; Inversion of spherically symmetric potentials from boundary data for the wave equation. Inverse Problems 14 (1998), no. 4, 999--1007.
 	\bibitem{Rakesh Webster's equation} Rakesh; Characterization of transmission data for Webster's horn equation. Inverse Problems 16 (2000), no. 2, L9--L24. 
 	\bibitem{Rakesh layered medium point source}Rakesh; An inverse problem for a layered medium with a point source.  Problems 19 (2003), no. 3, 497--506.
	\bibitem{Rakesh single-coincident source-receiver pair} Rakesh; Inverse problems for the wave equation with a single coincident source-receiver pair. Inverse Problems 24 (2008), no. 1, 015012, 16 pp.
	\bibitem{Rakesh and Sacks uniqueness angular controlled potential}Rakesh and P. Sacks; Uniqueness for a hyperbolic inverse problem with angular control on the coefficients. J. Inverse Ill-Posed Probl. 19 (2011), no. 1, 107--126. 
	\bibitem{Rakesh and Uhlmann inverse backscattering}Rakesh and G. Uhlmann; The point source inverse back-scattering problem. Analysis, complex geometry, and mathematical physics: in honor of Duong H. Phong, 279--289. 
		\bibitem{Romanov}V.G. Romanov;  On the problem of determining the coefficients in the lowest order terms of a hyperbolic equation. (Russian. Russian summary) Sibirsk. Mat. Zh. 33 (1992), no. 3, 156--160, 220; translation in Siberian Math. J. 33 (1992), no. 3, 497--500.
		\bibitem{Romanov problem of determining the two coefficients}V. G.  Romanov and D. I. Glushkova;  The problem of determining two coefficients of a hyperbolic equation. (Russian) Dokl. Akad. Nauk 390 (2003), no. 4, 452--456. 
		\bibitem{Romanov book integral geometry}V.G. Romanov;  Integral geometry and inverse problems for hyperbolic
	equations, volume 26. Springer Science and Business Media, 2013.
 	%\bibitem{Sacks and Symes uniqueness and continuous dependence}  P. Sacks  and W.W. Symes; Uniqueness and continuous dependence for a multidimensional hyperbolic inverse problem. Comm. Partial Differential Equations 10 (1985), no. 6, 635--676. 
	\bibitem{Santosa Syemes} F. Santosa and W.W. Symes; High-frequency perturbational analysis of the surface point-source response of a layered fluid. J. Comput. Phys. 74 (1988), no. 2, 318--381.	
%	\bibitem{Bukhgem and Klibanov uniqueness}Bukhge\~{i}m, A. L.; Klibanov, M. V. Uniqueness in the large of a class of multidimensional inverse problems. (Russian)
%	Dokl. Akad. Nauk SSSR 260 (1981), no. 2, 269--272.

	
	\bibitem{Sondhi} M.M. Sondhi;  A survey of the vocal tract inverse problem: theory, computations and experiments. Inverse problems of acoustic and elastic waves (Ithaca, N.Y., 1984), 1--19, SIAM, Philadelphia, PA, 1984.


	%\bibitem{Rakesh and Symes uniqueness result} Rakesh; Symes; Uniqueness for an inverse problem for the wave equation. Comm. Partial Differential Equations 13 (1988), no. 1, 87--96.

	\bibitem{Stefanov uniqueness}P.D.  Stefanov; A uniqueness result for the inverse back-scattering problem. Inverse Problems 6 (1990), no. 6, 1055--1064.
	
	

%	\bibitem{Klibanov partial point source} M.V. Klibanov; Some inverse problems with a `partial' point source. Inverse Problems 21 (2005), no. 4, 1379--1390.
%	\bibitem{Li partial point source} S. Li; Estimation of coefficients in a hyperbolic equation with impulsive inputs. J. Inverse Ill-Posed Probl. 14 (2006), no. 9, 891--904. 

	\bibitem{Symes} W. W. Symes; The seismic reflection inverse problem. Inverse Problems 25 (2009), no. 12, 123008, 39 pp.


	\bibitem{Manmohan} M. Vashisth; An inverse problems for the wave equation with source and receiver at distinct points, Journal of Inverse and Ill-posed Problems, http://doi.org/10.1515/jiip-2018-0004.
	
\end{thebibliography}

\documentclass[11pt,reqno]{amsart}
\usepackage[usenames]{color}
\usepackage{amsmath,verbatim,graphicx,epstopdf,enumerate}
\newcommand{\D}{\mathrm{d}}

\newcommand{\lb}{\left(}

\newcommand{\rb}{\right)}
\newcommand{\PD}{\partial}

\newcommand{\Rb}{\mathbb{R}}

\newcommand{\Beq}{\begin{equation}}
	\newcommand{\Eeq}{\end{equation}}
\newcommand{\beq}{\begin{equation*}}
	\newcommand{\eeq}{\end{equation*}}
\newcommand{\bal}{\begin{align}}
	\newcommand{\eal}{\end{align}}

\newcommand{\bp}{\begin{prob}}
	\newcommand{\ep}{\end{prob}}
\newcommand{\bpr}{\begin{proof}}
	\newcommand{\epr}{\end{proof}}

\newcommand{\bel}[1]{\begin{equation}\label{#1}}
	\newcommand{\ee}{\end{equation}}

\newcommand{\rr}{\mathbb{R}}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{prop}[theorem]{Proposition}

\theoremstyle{definition}

\title[An Inverse problem with under-determined data]{Unique determination of the damping coefficient in the wave equation using point source and receiver data}
\author[Vashisth]{Manmohan Vashisth}
\address{Beijing Computational Science Research Center, Beijing 100193, China.
	\newline
	\indent E-mail: {\tt manmohanvashisth@gmail.com}}

\begin{document}
\maketitle

\begin{abstract}
In this article, we consider the inverse problems of determining the damping coefficient appearing in the wave equation. We prove the unique determination of the coefficient  from the data coming from a single coincident source-receiver pair. Since our problem is under-determined, so some extra assumption on the coefficient is required to prove the uniqueness.
\end{abstract}\vspace{2mm}

\ \ \ \ \ \ \textbf{Keywords :} Inverse problems, wave equation, point source-receiver, damping coefficient\\\vspace{.5mm}

\ \ \ \ \ \ \textbf{Mathematics subject classification 2010:
 } 35L05, 35L10, 35R30, 74J25

\section{Introduction}
  We consider the following initial value problem (IVP),
\begin{equation}\label{wave equation with time derivative}
\begin{aligned}
(\Box- q(x)\partial_{t})u(x,t)&=\delta(x,t) \ \ \ \ (x,t)\in\rr^{3}\times\rr \\
u(x,t)|_{t<0}&=0\ \ \ \ \ \ \ \  \ \  \ \ \ \ \ \ \ \ \  x\in\rr^{3}
\end{aligned}
\end{equation}
where $\Box:=\PD_{t}^{2}-\Delta_{x}$ denotes the wave operator and the coefficient $q\in C^{\infty}(\Rb^{3})$ is known as damping coefficient. In this paper, we  study the problem of determination of coefficient $q$ appearing in \eqref{wave equation with time derivative} from the knowledge of solution measured at a single point for a certain period of time. We are interested in  the uniqueness of determination of coefficients $q$ from the knowledge of $u(0,t)$ for $t\in[0,T]$ with $T>0$ in Equation \eqref{wave equation with time derivative}. The problem studied here is motivated by geophysics, where geophysicists wish to determine the properties of earth structure by sending the waves from the surface of the earth and measuring the corresponding scattered responses (see \cite{Bube and Burridge 1-D,Symes} and references therein). Since the coefficient to be determined here depends on three variables while the given data depends on one variable as far as  the parameter count is concerned, the problem studied  here is under-determined. Thus some extra assumptions on coefficient $q$ are required in order to make the inverse problem solvable. We prove the uniqueness result for the radial coefficient.

There are several results related to the inverse problems for the wave equation with point source. We list them here. Romanov in \cite{Romanov} considered the problem for determining the  damping and potential coefficient in the wave equation with point source and proved unique determination of these coefficients by measuring the solution on a set containing infinite points. In \cite{Rakesh inversion of spherically symmetric potential} the problem of determining the radial potential from the knowledge of solution measured on a unit sphere for some time interval is studied. Rakesh and Sacks in \cite{Rakesh and Sacks uniqueness angular controlled potential}  established the  uniqueness for angular controlled potential in the wave equation from the knowledge of solution and its radial derivative measured on a unit sphere. In the above mentioned works the measurement set is an infinite set. Next we mention the work where uniqueness is established from the measurement of solution at a single point. Determination of the potential from the data coming from a single coincident source-receiver pair is considered in \cite{Rakesh single-coincident source-receiver pair} and the uniqueness result is established for the potentials which are either radial with respect a point different from source location or the potentials which are comparable. Recently author in \cite{Manmohan} extended the result of \cite{Rakesh single-coincident source-receiver pair} to a separated point source and receiver data. To  the best of our understanding,  very few results exist in the literature involving the recovery of the damping coefficient from point source and receiver data.  Our result, Theorem \ref{Damping Theorem 2}, is work in this direction. In the 1-dimensional inverse problems context, several results exist involving the uniqueness of recovery of the coefficient which depends on the space variable corresponding to  the first order derivative; see \cite{Ning_Yamamoto,Rakesh impedance inversion,Rakesh and Sacks impedance inversion,Rakesh Webster's equation,Romanov problem of determining the two coefficients,Sondhi}. We refer to \cite{Sacks_vocal_tract,Burridge integral equation,Levrentev Romanov and Shishatski book ill-posed,Rakesh layered medium point source,Rakesh and Uhlmann inverse backscattering} and references therein for more works related to the point source inverse problems for the wave equation.

We now state the main results of this article.

\begin{theorem}\label{Damping Theorem 2}
	Suppose $q_{i}(x)\in C^{\infty}(\rr^{3})$, $i=1,2$ with $q_{i}(x)=A_{i}(|x|)$ for some $C^{\infty}$ function $A_{i}$ on $[0,\infty)$. Let $u_{i}$ be the solution of the IVP
	\begin{align}\label{wave equation with time derivative2}
	\begin{aligned}
	(\Box- q_{i}(x)\partial_{t})u_{i}(x,t)&=\delta(x,t) \ \ \ \ (x,t)\in\rr^{3}\times\rr \\
	u_{i}(x,t)|_{t<0}&=0\ \ \ \ \ \ \ \  \ \  \ \quad \  x\in\rr^{3}.
	\end{aligned}
	\end{align}
	If $u_{1}(0,t)=u_{2}(0,t)$ for all $t\in[0,T]$ for some $T>0$, then $q_{1}(x)=q_{2}(x)$  for all $x$ with $|x|\leq T/2$, provided $q_{1}(0)=q_{2}(0)$.
\end{theorem}

The proof of the above theorem is based on an integral identity derived using the solution to an adjoint problem as used in \cite{Santosa Syemes} and \cite{Stefanov uniqueness}.  This idea was used in \cite{Eemeli,Rakesh and Uhlmann inverse backscattering,Manmohan} as well.

The article is organized as follows. In Section \ref{Prelim}, we state the existence and uniqueness results for the solution of  Equation \eqref{wave equation with time derivative},  the proof of which is given in \cite{Friedlander book,Levrentev Romanov and Shishatski book ill-posed,Romanov book integral geometry}.  Section \ref{Proof of Th 2} contains the proof of  Theorem \ref{Damping Theorem 2}.

\section{Preliminaries}\label{Prelim}
\begin{prop}\cite[pp.139,140]{Friedlander book}\label{Damping Prop}
	Suppose $q\in C^{\infty}(\Rb^{3})$
	and $u(x,t)$ satisfies the following initial value problem
	\begin{equation}\label{Hyperbolic PDE}
	\begin{aligned}
	Pu(x,t):=(\Box -q(x)\partial_{t})u(x,t)&= \delta(x,t),  \ (x,t)\in\rr^{3}\times\rr\\
	u(x,t)|_{t<0}\ &=\ 0, \quad \quad \quad \ \ \  x\in \rr^{3}
	\end{aligned}
	\end{equation}
	then $u(x,t)$ is given by 
	\begin{equation}\label{Fundamental solution to hyperbolic equation}
	u(x,t) =\frac{R(x,t)\delta(t-|x|)}{4\pi|x|} + v(x,t)
	\end{equation}
	where $v(x,t)=0$ for $t<|x|$ and in the region $t>|x|$, $v(x,t)$ is a $C^{\infty}$
	solution of the characteristic boundary value problem (Goursat Problem)
	\begin{equation}\label{Goursat problem}
	\begin{aligned}
	Pv(x,t)&= 0, \ for \ t>|x|\\
	v(x,|x|)&=-\frac{R(x,|x|)}{8\pi} \int\limits_{0}^{1}\frac{PR(sx,s|x|)}{R(sx,s|x|)} ds ,\ \ \  \forall x\in\rr^{3}
	\end{aligned}
	\end{equation}
and 	$R(x,t)$ is given by \cite[pp. 134]{Friedlander book}
	\begin{equation}\label{Expression for r}
	R(x,t)=\exp\left(-\frac{1}{2}\int\limits_{0}^{1}q(sx)t ds\right).
	\end{equation}
	
\end{prop}

\renewcommand{\thesection}{\large 3}
\section{Proof of Theorem \ref{Damping Theorem 2}}\label{Proof of Th 2}
In this section, we prove Theorem \ref{Damping Theorem 2}.   We will first prove an integral identity which will be used to prove our main result.
\begin{lemma}\label{Damping integral identity}
	Let $u_{i}(x,t)$ for $i=1,2$ be the solution to Equation \eqref{wave equation with time derivative2}. Then the following integral identity holds for all $\sigma\geq 0$
	\begin{equation}\label{Integral Identity for time derivative case}
	\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}u_{2}(x,t)u_{1}(x,2\sigma-t)dt dx =u(0,2\sigma)
	\end{equation}
	where $q(x):=q_{1}(x)-q_{2}(x)$ 
	and $u(x,t)=(u_{1}-u_{2})(x,t)$.\\
	\begin{proof} 
		Here we have  $u$ satisfies the following IVP 
		\begin{equation}\label{equation for u in time derivative case}
		\begin{aligned}
		\Box u(x,t)-q_{1}(x)\partial_{t}u(x,t)&=q(x)\partial_{t}u_{2}(x,t)\  \ \ \ \ (x,t)\in\rr^{3}\times\rr\\
		u(x,t)|_{t<0}&=0\ \ \ \ \ \ \ \ \ \ \  \ \ \ \ \ \ \ \ \ \ \ \ \ \ \  x\in\rr^{3}.
		\end{aligned}
		\end{equation}
		Multiplying Equation \eqref{equation for u in time derivative case} by $u_{1}(x,2\sigma-t)$ and integrating over $\rr^{3}\times\rr$, we have 
		\begin{align*}
		\begin{aligned}
	\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}u_{2}(x,t)u_{1}(x,2\sigma-t)dtdx &=	\int\limits_{\rr^{3}}\int\limits_{\rr}\left(\Box u(x,t) -q_{1}(x)\partial_{t}u(x,t)\right)u_{1}(x,2\sigma-t) dt dx\\
	&=\int\limits_{\Rb^{3}}\int\limits_{\Rb}u(x,t)\lb\Box u(x,t) -q_{1}(x)\partial_{t}u_{1}(x,2\sigma-t)\rb\D x \D t
		\end{aligned}
		\end{align*}
		where in the last step above we have used integration by parts and the properties of $v$ in Proposition \ref{Damping Prop}. Thus finally using the fact that $u_{1}$ is solution to \eqref{wave equation with time derivative2}, we get 
		\begin{align*}
		\begin{aligned}
		\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}u_{2}(x,t)u_{1}(x,2\sigma-t)dtdx=u(0,2\sigma);\ \mbox{for all $\sigma\geq 0$}.
		\end{aligned}
		\end{align*}
		 This completes the proof of the lemma.
	\end{proof}
\end{lemma}

Using Lemma \ref{Damping integral identity}   and  the fact that $u(0,t)=0$ for all $t\in[0,T]$, we see that
\begin{align*}
\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}u_{2}(x,t)u_{1}(x,2\sigma-t)dtdx=0; \mbox{ for all }\sigma\in[0,T/2].
\end{align*}
Now using Equation $\eqref{Fundamental solution to hyperbolic equation}$, we get 
\begin{align*}
\begin{aligned}
&\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}\Bigg(\frac{R_{2}(x,t)\delta(t-|x|)}{4\pi |x|}+v_{2}(x,t)\Bigg) \Bigg( \frac{R_{1}(x,2\sigma-t)\delta(2\sigma-t-|x|)}{4\pi |x|}\Bigg)dtdx\\
&+ \int\limits_{\rr^{3}}\int\limits_{\rr}q(x)v_{1}(x,2\sigma-t)\partial_{t}\Bigg(\frac{R_{2}(x,t)\delta(t-|x|)}{4\pi |x|}+v_{2}(x,t)\Bigg)dtdx=0.
\end{aligned}
\end{align*}
This gives
\begin{equation}\label{Explicit form of integral identity in time derivative case}
\begin{aligned}
&\underbrace{\int\limits_{\rr^{3}}\int\limits_{\rr} \frac{q(x)\partial_{t}R_{2}(x,t)R_{1}(x,2\sigma-t)\delta(t-|x|)\delta(2\sigma-t-|x|)}{16\pi^{2}|x|^{2}}dtdx}_{I_{1}}\\ &+\underbrace{\int\limits_{\rr^{3}}\int\limits_{\rr}\frac{q(x)R_{2}(x,t)R_{1}(x,2\sigma-t)\partial_{t}\delta(t-|x|)\delta(2\sigma-t-|x|)}{16\pi^{2}|x|^{2}}dtdx}_{I_{2}}\\
&+\underbrace{\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}\Big(\frac{R_{2}(x,t)\delta(t-|x|)}{4\pi |x|}\Big)v_{1}(x,2\sigma-t)dtdx}_{I_{3}}\\
&+\underbrace{\int\limits_{\rr^{3}}\int\limits_{\rr}\frac{q(x)\partial_{t}v_{2}(x,t)R_{1}(x,2\sigma-t)\delta(2\sigma-t-|x|)}{4\pi |x|}dtdx}_{I_{4}}\\
&+\underbrace{\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}v_{2}(x,t)v_{1}(x,2\sigma-t)dtdx}_{I_{5}}=0;\mbox{ for all }  \sigma\in [0,T/2].
\end{aligned}
\end{equation}
In a compact form, this can be written as
\begin{equation}\label{integral identity in terms of Ij}
I_{1}+I_{2}+I_{3}+I_{4}+I_{5}=0.
\end{equation}
Next we simplify each $I_{j}$ with $j=1,2,.....,5$. We will use the fact that $v_{i}(x,t)=0$ for $t<|x|$. 

We have 
\begin{align*}
I_{1}&=\int\limits_{\rr^{3}}\int\limits_{\rr} \frac{q(x)\partial_{t}R_{2}(x,t)R_{1}(x,2\sigma-t)\delta(t-|x|)\delta(2\sigma-t-|x|)}{16\pi^{2}|x|^{2}}dt dx\\
&=\int\limits_{|x|=\sigma}\frac{q(x)\partial_{t}R_{2}(x,|x|)R_{1}(x,|x|)}{16\pi^{2}|x|^{2}}dS_{x}\\
&=-\int\limits_{|x|=\sigma}\frac{q(x)R_{1}(x,|x|)R_{2}(x,|x|)}{32\pi^{2}|x|^{2}}\Big(\int\limits_{0}^{1}q_{2}(sx)ds\Big)dS_{x}.
\end{align*}
Next we simplify the integral $I_{2}$. We use the following formula \cite[Page 231, Eq.(10)]{Gelfand}
\begin{equation}\label{Dirac derivative}
\int \delta'(r-|x|)\varphi dx=\frac{-1}{|x|^{2}}\int\limits_{|x|=r} \frac{\partial }{\partial r} \left (\varphi r^{2}\right) dS_{x}.
\end{equation}
Note that from  this formula, by a change of variable, we have
\begin{equation}\label{Dirac derivative 2}
\int \delta'(2r-2|x|)\varphi dx=\frac{-1}{2|x|^{2}}\int\limits_{|x|=r} \frac{\partial }{\partial r} \left (\varphi r^{2}\right) dS_{x}.
\end{equation}
Now
\begin{align*}
I_{2}&=\int\limits_{\rr^{3}}\int\limits_{\rr}\frac{q(x)R_{2}(x,t)R_{1}(x,2\sigma-t)\partial_{t}\delta(t-|x|)\delta(2\sigma-t-|x|)}{16\pi^{2}|x|^{2}}dtdx\\
&=\int\limits_{\rr^{3}}\int\limits_{\rr}\frac{q(x)R_{2}(x,t)R_{1}(x,2\sigma-t)\delta^{'}(t-|x|)\delta(2\sigma-t-|x|)}{16\pi^{2}|x|^{2}}dtdx\\
&=\int\limits_{\rr^{3}}\frac{q(x)R_{2}(x,2\sigma-|x|)R_{1}(x,|x|)\delta'(2\sigma-2|x|)}{16\pi^{2}|x|^{2}}dx\\
&=-\frac{1}{32\pi^{2}\sigma^{2}}\int\limits_{|x|=\sigma}\frac{\partial}{\partial{r}}\{q(x)R_{1}(x,|x|)R_{2}(x,2\sigma-|x|)\}dS_{x}.
\end{align*}
In the last step above, we used Equation \eqref{Dirac derivative 2}.

Next we have 
\begin{align*}
I_{3}&=\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}\Big(\frac{R_{2}(x,t)\delta(t-|x|)}{4\pi |x|}\Big)v_{1}(x,2\sigma-t)dxdt.\\
\intertext{{We can view the derivative above as a limit of the difference quotients in the distribution topolgy \cite[pp.48]{Friedlander-Distributions-book}. Combining this with the fact that $v_{1}$ is $C^{2}$ in $\{(x,t):|x|\leq t\}$, we get,}}
I_{3}&=-\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\frac{R_{2}(x,t)\delta(t-|x|)}{4\pi |x|}\partial_{t}\Big(v_{1}(x,2\sigma-t)\Big)dxdt\\
&=\int\limits_{\rr^{3}}\frac{q(x)R_{2}(x,|x|)\partial_{t}v_{1}(x,2\sigma-|x|)}{4\pi |x|}dx.\\
\intertext{Again using the fact that $v_{1}(x,t)=0$ for $t<|x|$, we get,}
&I_{3}=\int\limits_{|x|\leq \sigma}\frac{q(x)R_{2}(x,|x|)\partial_{t}v_{1}(x,2\sigma-|x|)}{4\pi |x|}dx.
\end{align*}
Next we simplify $I_{4}$. Similiar to $I_{3}$, we have
\begin{align*}
I_{4}&=\int\limits_{\rr^{3}}\int\limits_{\rr}\frac{q(x)\partial_{t}v_{2}(x,t)R_{1}(x,2\sigma-t)\delta(2\sigma-t-|x|)}{4\pi |x|}dtdx\\
&=\int\limits_{|x|\leq \sigma}\frac{q(x)R_{1}(x,|x|)\partial_{t}v_{2}(x,2\sigma-|x|)}{4\pi|x|}dx.
\end{align*}
Finally, we have
\begin{align*}
I_{5}&=\int\limits_{\rr^{3}}\int\limits_{\rr}q(x)\partial_{t}v_{2}(x,t)v_{1}(x,2\sigma-t)dtdx\\
&=\int\limits_{|x|\leq\sigma}\int\limits_{|x|}^{2\sigma-|x|}q(x)\partial_{t}v_{2}(x,t)v_{1}(x,2\sigma-t)dtdx.
\end{align*}
Now, we use the fact that $q_i$ is a radial function, that is, $q_{i}(x)=A_{i}(|x|)$.  Then note that
\[
R_{i}(x,|x|)= \exp\left(-\frac{|x|}{2} \int\limits_{0}^{1} q_{i}(sx)ds\right)=\exp\left(-\frac{|x|}{2} \int\limits_{0}^{1} A_{i}(s|x|)ds\right)
\]
is also radial. For simplicity, we denote $R(x,|x|)$ by $R(|x|)$.

With this, we have
\[
I_{1}=-\frac{A(\sigma)R_{1}(\sigma)R_{2}(\sigma)}{8\pi}\int\limits_{0}^{1} A_{2}(s\sigma) ds.
\]
Next we consider $I_{2}$. First let us consider the derivative:
\[
D_{r}:=\frac{\partial}{\partial r}\left( A(r) R_{1}(x,r) R_{2}(x,2\sigma-r)\right).
\]
After a routine calculation, we get,
\begin{align*}
D_{r}&= A'(r) R_{1}(x,r)R_{2}(x,r) -\frac{1}{2} A(r)^{2} R_{1}(x,r)R_{2}(x,2\sigma-r) \\
&-\sigma A(r)R_{1}(x,r)R_{2}(x,2\sigma-r) \int\limits_{0}^{1} A_{2}'(rs)s ds\\
&= A'(r) R_{1}(x,r)R_{2}(x,r) -\frac{1}{2} A(r)^{2} R_{1}(x,r)R_{2}(x,2\sigma-r) \\
&-A(r)R_{1}(x,r)R_{2}(x,2\sigma-r)\left[\frac{\sigma}{r}\left(A_{2}(r)-\int\limits_{0}^{1}A_{2}(rs) ds\right)\right].
\end{align*}
On $|x|=\sigma$, we have
\begin{align*}
D_{r}|_{|x|=\sigma}&=R_{1}(\sigma)R_{2}(\sigma)\left[A'(\sigma)-\frac{1}{2} A(\sigma)^{2}-A(\sigma)A_{2}(\sigma)+A(\sigma)\int\limits_{0}^{1}A_{2}(s\sigma) ds\right]\\
&=R_{1}(\sigma)R_{2}(\sigma)\left[A'(\sigma) -\frac{1}{2}A(\sigma)(A_{1}+A_{2})(\sigma)+A(\sigma)\int\limits_{0}^{1}A_{2}(s\sigma) ds\right].
\end{align*}
Hence
\begin{align*}
I_{2}&=-\frac{1}{8\pi}\left(R_{1}(\sigma)R_{2}(\sigma)\left[A'(\sigma) -\frac{1}{2}A(\sigma)(A_{1}+A_{2})(\sigma)+A(\sigma)\int\limits_{0}^{1}A_{2}(s\sigma) ds\right]\right).
\end{align*}
Let us denote 
\[
\tilde{A}(\sigma)=A(\sigma)R_{1}(\sigma)R_{2}(\sigma).
\]
Then
\begin{align*}
I_{2}=-\frac{1}{8\pi} \frac{d}{d\sigma}\tilde{A}(\sigma)-\frac{1}{8\pi}\tilde{A}(\sigma)\int\limits_{0}^{1}A_{2}(s\sigma)ds.
\end{align*}
Therefore
\[
I_{1}+I_{2}=-\frac{1}{8\pi} \left(2\tilde{A}(\sigma)\int\limits_{0}^{1}A_{2}(s\sigma) ds+\frac{d}{d\sigma}\tilde{A}(\sigma)\right).
\]
Considering the following integrating factor for $I_{1}+I_{2}$ 
\[
\exp\left(2\int\limits_{0}^{\sigma}\int\limits_{0}^{1}A_{2}(ts)dt ds\right),
\]
we have
\[
I_{1}+I_{2}=-\frac{1}{8\pi}\exp\left(-2\int\limits_{0}^{\sigma}\int\limits_{0}^{1}A_{2}(ts)dt ds\right)\frac{d}{d\sigma}\left[\exp\left(2\int\limits_{0}^{\sigma}\int\limits_{0}^{1}A_{2}(ts)dt ds\right)\tilde{A}(\sigma)\right].
\]

Now from Equation \eqref{integral identity in terms of Ij},     we have
\begin{equation}\label{integral equation in radial case}
\begin{aligned}
& \frac{1}{8\pi}\frac{d}{d\sigma}\left[\tilde{A}(\sigma)\exp\left(2\int\limits_{0}^{\sigma}\int\limits_{0}^{1}A_{2}(st)dsdt\right)\right]\\
&\ \ =\exp\left(2\int\limits_{0}^{\sigma}\int\limits_{0}^{1}A_{2}(st)dsdt\right)\Bigg[\int\limits_{|x|\leq \sigma}\frac{q(x)R_{2}(x,|x|)\partial_{t}\{R_{1}v_{1}\}(x,2\sigma-|x|)}{4\pi |x|}dx\\
&\ \ \ \ +\int\limits_{|x|\leq\sigma}\frac{q(x)R_{1}(x,|x|)\partial_{t}v_{2}(x,2\sigma-|x|)}{4\pi|x|}dx\\
&\ \ \ \ +\int\limits_{|x|\leq\sigma}\int\limits_{|x|}^{2\sigma-|x|}q(x)\partial_{t}v_{2}(x,t)v_{1}(x,2\sigma-t)dtdx\Bigg]\mbox{ for all } \sigma\in[0,T/2].
\end{aligned}
\end{equation}
Integrating on both sides with respect to $\sigma$ under the assumption  that $\tilde{A}(0)=0$, we get 
\begin{align*}
&\exp\left(\int\limits_{0}^{\tilde{\sigma}}\int\limits_{0}^{1}2 A_{2}(st)dsdt\right)\tilde{A}(\tilde{\sigma})\\
&=\int\limits_{0}^{\tilde{\sigma}}\exp\left(\int\limits_{0}^{\sigma}\int\limits_{0}^{1}2 A_{2}(st)dsdt\right)\Bigg\{\int\limits_{|x|\leq \sigma}\frac{q(x)R_{2}(x,|x|)\partial_{t}v_{1}(x,2\sigma-|x|)}{4\pi |x|}dx\\
&+\int\limits_{|x|\leq\sigma}\frac{q(x)R_{1}(x,|x|)\partial_{t}v_{2}(x,2\sigma-|x|)}{4\pi|x|}dx\\
&+\int\limits_{|x|\leq\sigma}\int\limits_{|x|}^{2\sigma-|x|}q(x)\partial_{t}v_{2}(x,t)v_{1}(x,2\sigma-t)dtdx\Bigg\} d\sigma, \  \mbox{ for all } \tilde{\sigma}\in[0,T/2].
\end{align*}
Now using the fact that $R_{i}'s$ are continuous, non-zero functions, and $v_{i}'s$ are continuous, we have the following inequality:
\begin{align*}
|\tilde{A}(\tilde{\sigma})|\leq C\int\limits_{0}^{\tilde{\sigma}}|\tilde{A}(r)| dr\mbox{ for all }\tilde{\sigma}\in[0,T/2].
\end{align*}
Now by Gronwall's inequality, we have $\tilde{A}(\sigma)=0$ for all $\tilde{\sigma}\in [0,T/2]$, which gives us $q_{1}(x)=q_{2}(x)$ for all $x\in\rr^{3}$ such that $|x|\leq T/2$. This completes the proof.

\section*{Acknowledgement}
The author would like to thank Dr. Venky Krishnan for useful discussions. He is supported by NSAF grant (No. U1530401).

\end{document}